2x²+ (1/3)x²-6=5
x=

Answer by ewatrrr(24785)

Previously Posted
2x²+ (1/3)x²-6=5
(7/3)x^2 = 11
x^2 = (3/7)11 = 33/7
x = ± √(33/7)